# ChIRAAG: ChatGPT Informed Rapid and Automated Assertion Generation

### v2

@@ Abstract
System Verilog Assertion (SVA) formulation- a critical yet complex task is a prerequisite in the Formal Property Verification (FPV) process. Traditionally, SVA formulation involves expert-driven interpretation of specifications, which is time-consuming and prone to human error. However, LLM-informed automatic assertion generation is gaining interest. We designed a novel framework called ChIRAAG, based on OpenAI GPT4, to generate SVA assertions from natural language specifications. ChIRAAG constitutes the systematic breakdown of design specifications into a standardized format, further generating assertions from formatted specifications using LLM. Furthermore, we developed testbenches to verify/validate the LLM-generated assertions. Automatic feedback of log files from the simulation tool to the LLM ensures that the framework can generate correct SVAs automatically. Only 33% of LLM-generated raw assertions had errors. Our results on OpenTitan designs shows that LLMs can streamline and assist engineers in the assertion generation process, reshaping verification workflows.
@@ I Introduction
Large Language Models (LLMs) have evolved to assist humans in solving complex problems. LLMs create numerous possibilities for transformative impact across different domains, including education [1], medicine [2], and finance [3]. The hardware design industry has used LLMs to execute steps of Electronic Design Automation (EDA) at a faster rate[4].
Pre-silicon bug detection is a crucial step in designing functionally correct hardware. Security vulnerabilities and bugs are identified during the pre-silicon phase via Formal Property Verification (FPV). As per the functional verification study conducted by Wilson Research Group in 2020 [5], over 50% of the overall development resources and expenditures in Application-Specific Integrated Circuit (ASIC) and Field-Programmable Gate Array (FPGA) systems were allocated to FPV. FPV is a systematic and rigorous process of using formal methods to verify whether a system adheres to its specifications. FPV ensures that the system meets specified properties under all possible scenarios [6].
Assertions Based Verification (ABV) are used to verify the properties in over 50% of the designs  [5]. ABV validates the correct design implementation by integrating safety and liveness properties, expressed as assertions, throughout the development process. Safety property states that any path (sequence of events or states) violating the property has a finite prefix (initial segment of the path) such that every extension of the prefix violates the property. For example, “whenever a signal request is asserted, signal acknowledgment is asserted within three cycles” is a safety property [6]. While a liveness property states that any finite path (sequence of events or states) can be extended to a path satisfying the property [6]. For example, “Every request is eventually acknowledged” is a liveness property. However, as the scale of design and verification requirements increases, so does the number of assertions, which can reach hundreds, making it too difficult to be proven formally or falsified within the constraints of available resources manually [7].
LLMs could perform ABV quickly by automatically generating the correct assertions from the natural language specification that holds in simulation. The existing methods have relied on generating SVA from Verilog code [8], comments, and assertion-based examples [9]. None of them could generate SVA assertions directly from specification. Also, they still require potential intervention and expertise from the experts. This is the motivation for this study to formulate a novel framework that reduces human effort and time in ABV. This work aims to analyze how LLMs can assist ABV by generating assertions automatically from the natural language specifications. In our framework, manual intervention is needed only when there is a potential bug in the design implementation. We seek to answer two key questions through this work: RQ1: How good are LLMs in generating System Verilog Assertion (SVA) from natural language specification? RQ2: How functionally correct are the generated assertions within the design and verification framework?.
@@
  • A prompting method to provide relevant information to an LLM to generate SVA.
@@ II Background and related works
II-A Overview of LLM
LLMs are Transformer Language models where the construction of the weight parameters involves training models using an extensive volume of text data acquired from diverse sources such as various websites, web-books, and PDFs. Based on their functionality, LLMs can be divided into two: General Domain LLMs (gd-LLMs) and Domain-specific LLMs (ds-LLMs) [11]. gd-LLMs are trained on diverse resources from various domains and sources, making them applicable to a wide range of queries. Examples of gd-LLMs include Llama 2 [12], and GPT4 [13]. While ds-LLMs focus on a particular realm of interest. Examples include MedicalGPT [14] and FoodGPT [15]. LLMs are expensive to train on large datasets, so fine-tuning pre-trained LLMs is becoming important to simplify training. Fine-tuning of an LLM is a supervised process to adjust the model parameters based on the task [16]. Code Llama [17] by Meta and StarCoder [18] by Hugging Face are some examples.
II-B LLM-Driven Automatic Assertion Generation
@@
LLMs have leading-edge capabilities in diverse tasks, including source code generation [19], automated testing [20], and program repair [21, 22]. Automatic assertion generation is gaining popularity in recent times [8]. This study showcased the capability of LLM to generate assertions solely from RTL code without additional specifications. Despite initial syntactic and semantic errors, the authors observed that, with proper guidance, LLMs consistently produce correct and complete assertions. The work in [9] uses comments about the security assertions to generate assertions. They assessed the effectiveness of LLMs in accelerating property verification.They generated 75,600 assertions across diverse conditions and varying levels of context and found that LLMs can generate correct assertions at an average rate of 4.53%. They also mentioned that detailed comments and context proved to be particularly effective in generating correct assertions. The work in [23] uses LLMs in security assertion generation. They perform formal property verification to identify vulnerabilities before developing the Design Under Test (DUT). The generated assertions at the pre-silicon level can help monitor and mend security vulnerabilities during the initial phases of design.
@@
In contrast, we have exclusively used the design specification in natural language as an input to LLM to generate SV assertion. The design specification operates at an abstract level, serving as a representation of the design, in contrast to directly providing RTL code for assertion generation. Since the objective of assertions is to identify the bug in the implementation. Therefore, it is not desirable to obtain the assertions from implementation. Moreover, creating assertions from a design specification poses a significant challenge in the absence of a dedicated and specialized framework. Therefore, we proposed a novel framework that formats the input specification to the LLM and further reviews, analyzes, and mends the LLM-generated assertions, if required. We have mainly focused on generating functional assertions rather than generating security assertions. None of the above-discussed works generate assertions merely using specification.
II-C The Three C’s in ABV
ABV uses assertions, which are formal statements or conditions that must always hold in a design for the proper functioning of logic. A primary objective of an assertion is to ensure consistency in the context of the designer’s intent and the actual realization of the design [24]. “Consistency” denotes no conflict or contradiction between the formal properties/specifications that could lead to ambiguous or erroneous results. “Correctness” refers to the properties where it assures that the assertions correctly represent the intended behavior of the design. “Completeness” states that the assertions should cover all related conditions and edge cases within the design. The verification process shouldn’t miss critical aspects of system behavior. For example, in a simple digital counter,
assert @(posedge clk_i) (count == previous_count + 1);
represent a correct assertion that ensures the counter increments correctly with each clock cycle by checking that the current count equals the previous count plus one. While,
assert @(posedge clk_i) (enable_signal == (start_signal && !stop_signal));
represents a consistent assertion that ensures control signals are handled consistently. This case checks that the enable signal is true only when the start signal is true and the stop signal is false, preventing contradictory combinations of control signals in the digital counter circuit. And,
assert @(posedge clk_i) (count >= 0 && count <= MAX_COUNT);
represents a completeness assertion that ensures that the counter’s value remains within the valid range. It checks that the count is always $\geq 0$ and does not exceed the maximum count value (MAX_COUNT). The main motivation of our work is to assist the engineer in generating the assertions using LLMs directly from the specifications. In this work, we mainly focused on the correctness of the assertions that are generated by the LLMs.
@@ III ChIRAAG: Our SVA Generation Framework
The automatic assertion generation with ChIRAAG involves the compilation of two significant stages. Algorithm 1 shows the steps of our proposed ChIRAAG framework. While the stages are briefly described below:
@@ III-A Formatting of specification
The formatting of design specifications includes representing the design context with specific labels informing about its general definitions, signals, functional parameters, etc. The formatting involves extracting the important information from the design specification, $\mathcal{D}$, and later converting the extracted information, $UnFmt$, into a Javascript Object Notation (JSON) format, $\mathcal{J}$. This operation can be symbolically represented as $Extraction:\mathcal{D}\rightarrow UnFmt$ followed by $JSON:UnFmt\rightarrow\mathcal{J}$, as represented in Line 7-10 of Algorithm 1. The $JSON$ function comprises several sub-functions, each responsible for structuring a specific aspect of the extracted information into JSON format, $\mathcal{J}$, where $\mathcal{J}$ includes {{\$\mathcal{J}_{int}$, $\mathcal{J}_{s\_ov}$, $\mathcal{J}_{df}$, $\mathcal{J}_{par}$, $\mathcal{J}_{fn\_r}$, $\mathcal{J}_{tm\_r}$, $\mathcal{J}_{ext}$}}\}}. The components in Table I show the formatting functions with their respective labels.
@@ III-B Automatic Assertion Generation
This is the second stage of ChIRAAG, which mainly involves the generation of the SV assertions using the JSON formatted specifications. The formatted specifications of the designs are provided as input to the LLM through our framework by making a call to OpenAI Application Programming Interface (API). This generates raw SV assertions as stated in Line 13 of Algorithm 1. The generated assertions are further verified for correctness through simulation runs utilizing testbenches in a simulation environment, producing simulation logs. This step gives rise to four possible cases that can eventually occur. They are:
No Error Message
@@
When the generated assertion, $A_{n}$, passes all the test cases, i.e., the log file, $\mathcal{L}$, contains no errors, then $A_{n}$ is taken as correct output assertions suite, $S$. Lines 15-16 of Algorithm 1 shows that when $\mathcal{L}$ contains no error message, the while loop gets terminated, and then the ‘OUTPUT GENERATION’ procedure, in Lines 34-37, gets executed.
Syntax Error
@@
Some LLM-generated assertions have Syntax Errors (SE). LLM is automatically instructed about the syntax error and also to refine the assertion. It is able to correct the syntax error within a few iterations. Lines 17-18 of Algorithm 1 check if the log file, $\mathcal{L}$, contains any SE, then the function ‘SyntaxRectifier’ gets executed. ‘SyntaxRectifier’ takes the assertions, $A_{n}$, and log file, $\mathcal{L}$, as input, and a syntax-based prompt as a heading is provided automatically to the LLM using the file-handling process. This process generates the refined assertions, $\overline{A_{n}}$.
Simulation Failure
The assertions generated by LLM might fail in simulation run because of specific errors. Two such simulation errors we have encountered are timing errors and missing signals. Lines 20-21 of Algorithm 1 check if the log file, $\mathcal{L}$, has any timing or missing signal error, then similar to SE, here the function ‘SimulationRectifier’ takes the assertions, $A_{n}$, and the log file, $\mathcal{L}$, as input and further a simulation-based prompt as a heading is provided automatically to LLM using the file-handling process. This leads to the generation of the rectified assertions, $\overline{A_{n}}$.
Implementation Bug
@@
Sometimes, it might be possible that even the generated assertions are correct, but the implementation might be buggy. In such a scenario, the testbench may give out a wrong output, resulting in testcase failure. In such cases, manual inspection of design implementation is to be done as shown in Lines 23-24 of Algorithm 1.
@@
The assertions may not pass the test cases in a single run. Therefore, we iteratively run the error-checking process multiple times. Line 14 of Algorithm 1 shows that the while loop runs until $n$ is less than the upper bound $\mathcal{T}$. If $n$ gets equal to $\mathcal{T}$, and there are still error messages in the log file, $\mathcal{L}$, we manually examine the design implementation for potential bugs, as stated in Lines 29-33. Also, the newly generated assertions, $\overline{A_{n}}$, is asserted to $A_{n}$, as stated in Lines 19 and 22 of the Algorithm 1, to follow the further error-checking execution steps in the loop. If there will be no error message in the log file, $\mathcal{L}$, and $n$ is less than $\mathcal{T}$, we execute the Lines 34-37 of Algorithm 1 as discussed in `$No\text{ }Error\text{ }Message^{\prime}$ case to get our final SV assertions, $S$. The proposed ChIRAAG framework is illustrated in the Fig. 1.
@@
Algorithm 1 ChIRAAG: Automated SVA Generation Framework
@@
2:      $\mathcal{D}$: Design specification document
3:      $\mathcal{L}$: Log file of simulation tool.
4:      $\mathcal{T}$: Upper bound on the number of LLM interaction prompts
5:Output:
6:      $S$: Suite of System Verilog Assertions
7:procedure Specification Extraction and Formatting
8:    $UnFmt\leftarrow Extraction(\mathcal{D})$
9:    $\mathcal{J}\leftarrow\text{JSON}(UnFmt)$
10:end procedure
11:procedure LLM Interaction for SVA Generation
12:    Initialize $n\leftarrow 0\text{ and }\mathcal{L}\leftarrow\{\}$
13:    $A_{n}\leftarrow\text{LLM}(\mathcal{J},\mathcal{L})$
14:    while $n<\mathcal{T}$ do
15:         if no errors in  of $\mathcal{L}\text{ of }A_{n}$ then
16:             break
17:         else if syntactic errors in  of $\mathcal{L}\text{ of }A_{n}$ then
18:             $\overline{A}_{n}\leftarrow\text{SyntaxRectifier}(A_{n},\mathcal{L})$
19:             $A_{n}\leftarrow\overline{A_{n}}$
20:         else if timing or missing signal error in  of $\mathcal{L}\text{ of }A_{n}$ then
21:             $\overline{A}_{n}\leftarrow\text{SimulationRectifier}(A_{n},\mathcal{L})$
22:             $A_{n}\leftarrow\overline{A_{n}}$
23:         else if wrong output due to testcase failure then
24:             Manual inspection of design implementation
25:             break
26:         end if
27:         $n\leftarrow n+1$
28:    end while
29:end procedure
30:procedure Design Implementation Examination
31:    if $n==\mathcal{T}$ and error(s) in $\mathcal{L}$ then
32:         Examine design implementation for bugs
33:    end if
34:end procedure
35:procedure Output Generation
36:    $S\leftarrow{A}_{n}$
37:    return $S$
38:end procedure
@@ IV Case Study: RV Timer
Our proposed Algorithm 1 on SVA generation is assessed using multiple designs. However, we are taking a particular design, RV Timer, to mention our findings.
The RV Timer has clk_fixed as the fixed clock input and prescaler that divides the input clock frequency to a lower frequency. This lower frequency is later used by the timer. The timer’s count is incremented based on timer ticks generated, which is determined by the prescaler value. The generated ticks are sent to multiple hardware threads. A step value serves as an input to the timer, determining the increment step for each tick. The mtime register is the main timer counter that gets incremented with each tick and holds the current tick value. The comparison to check if a timer has reached a specific time value is performed using the mtimecmp register. The output signal expired is triggered when the mtime value is greater than or equal to the mtimecmp register value. The block diagram of the RV Timer is shown in Fig . 2
@@
The unstructured design specification of RV Timer is initially transformed to structured JSON format and presented as input to the LLM to generate the SV assertions. Listing 1 provides a part of the structured format specification of RV Timer.
Listing 1: RV Timer Module Specification
"Introduction": {
"Purpose": "This document specifies the operational, functional, and timing requirements for the RV Timer module, which is designed to provide timing functions for the system.",
"System_Overview": {
"Architecture_Summary": "The RV Timer module includes tick generators, counters, comparators, and is capable of raising interrupts based on configurable timing events.",
"Definitions": {
"Signals": [
"Name": "clk_fixed",
"Type": "Input",
"Width": "As per system configuration",
"Description": "Fixed clock input for the timer."
"Parameters": [
"Name": "prescaler",
"Type": "Configuration",
"Description": "Prescaler value for the fixed clock to generate the timer tick."
"Functional_Requirements": {
The LLM initially generated 8 functional assertions. We further evaluated the correctness of the raw assertions using our testbench and simulator. Of the 8 assertions generated, 5 were correct in their raw format. This shows the potential of our proposed framework ChIRAAG. The remaining 3 raw assertions were rectified automatically using the steps of Algorithm 1. This process generated an extra assertion, which is essential to cover all the design aspects, increasing the count of generated assertions to 9. Some of the initial assertions of RV Timer generated by the LLM are:
@@
The Assertion 1 is a correct raw assertion that passes all the tests. It ensures that ‘tick_count’ resets to 0 when reset is asserted, i.e., when ‘rst_ni’ or the active low reset signal is not active. While the majority of assertions are correct in their raw format, a few failed to pass certain test cases due to the presence of syntactic error or simulation error. As in Assertion 2, we observed that a missing signal prevents passing all the test cases as shown in Fig. 3. The LLM understood the missing signal rst_ni from the learned information and included it in the generated assertion. The ‘SimulationRectifier’ function shown in Algorithm 1 is called upon to rectify the Assertion 2. The LLM generated the refined assertion and modified it into two distinct assertions Modified Assertion 2a and Modified Assertion 2b.
Figure 3: Failure log of Missing signal
@@
The Modified Assertion 2a ensures that as long as the system is not in reset, is active, and the tick count is below a set threshold, the tick count will increment by 1 in each clock cycle. The Modified Assertion 2b verifies that when the system transitions from a reset state to a non-reset state, the tick count is reset to 0, ensuring proper initialization or re-initialization of the system’s tick counter. Together, Modified Assertion 2a and Modified Assertion 2b form a comprehensive set of checks for the tick_count signal, covering both its incremental behavior during normal operation and its initialization behavior upon reset deassertion. These assertions ensure the tick_count signal behaves correctly across different operational states of the system, contributing to the overall reliability of the hardware functioning.

### v3

@@ Abstract
System Verilog Assertion (SVA) formulation- a critical yet complex task is a prerequisite in the Assertion Based Verification (ABV) process. Traditionally, SVA formulation involves expert-driven interpretation of specifications, which is time-consuming and prone to human error. Recently, LLM-informed automatic assertion generation is gaining interest. We designed a novel framework called ChIRAAG, based on OpenAI GPT4, to generate SVA from natural language specifications of a design. ChIRAAG constitutes the systematic breakdown of design specifications into a standardized format, further generating assertions from formatted specifications using LLM. Furthermore, we used few test cases to validate the LLM-generated assertions. Automatic feedback of log messages from the simulation tool to the LLM ensures that the framework can generate correct SVAs. In our experiments, only 27% of LLM-generated raw assertions had errors, which was rectified in few iterations based on the simulation log. Our results on OpenTitan designs show that LLMs can streamline and assist engineers in the assertion generation process, reshaping verification workflows.
@@ I Introduction
Pre-silicon bug detection is a crucial step in designing functionally correct hardware. As per the functional verification study conducted by Wilson Research Group in 2020 [1], over 50% of the overall development resources and expenditures in Application-Specific Integrated Circuit (ASIC) and Field-Programmable Gate Array (FPGA) systems were allocated to the verification process. Assertion Based Verification (ABV) validates the correct design implementation by integrating safety and liveness properties, expressed as assertions, throughout the development process. Safety property states “something bad does not happen”. While liveness property states “something good will eventually happen” [2]. However, as the scale of design and verification requirements increases, so does the number of assertions, which can reach hundreds. Identifying and writing manually all assertions of a design is a non-trivial task.
Large Language Models (LLMs) have evolved to assist humans in solving complex problems. LLMs can also help in ABV by automatically generating the assertions from the natural language specification of a design. The existing methods have relied on generating SVA from Verilog code [3], comments, and assertion-based examples [4]. None of them could generate SVA directly from the specifications. This is the motivation for this study to formulate a novel framework that reduces human effort and time in SVA formulation. This work aims to analyze how LLMs can assist SVA generation semi-automatically from the natural language specifications. In our framework, manual intervention is needed when there is a potential bug in the implementation or the LLM cannot fix the issue in the generated assertion(s) based on the simulation log. We seek to answer two key questions through this work: RQ1: How good are LLMs in generating System Verilog Assertion (SVA) from natural language specification? RQ2: How functionally correct are the generated assertions within the design and verification framework?.
@@
  • The standardization of the format for specification and a prompting method to generate SVA.
@@
The rest of the paper is organized as follows: Section \@slowromancapii@ discusses the related works. Section \@slowromancapiii@ describes our proposed framework. Section \@slowromancapiv@ shows a case study on a particular design. The experimentation and performance results of our proposed approach are illustrated in Section \@slowromancapv@. Section \@slowromancapvi@ concludes the paper with future directions.
@@ II Background and related works
LLMs are Transformer Language models where the construction of the weight parameters involves training models using an extensive volume of text data acquired from diverse sources such as various websites, web-books, and PDFs. LLMs have leading-edge capabilities in diverse tasks, including source code generation [6], automated testing [7], and program repair [8]. As in [3], the authors showcased the capability of LLM to generate assertions solely from RTL code without additional specifications. Despite initial syntactic and semantic errors, the authors observed that, with proper guidance, LLMs consistently produce correct and complete assertions. The work in [4] uses comments about the security assertions to generate assertions. They assessed the effectiveness of LLMs in accelerating property verification. They generated 75,600 assertions across diverse conditions and varying levels of context and found that LLMs can generate correct assertions at an average rate of 4.53%. The work in [9] uses LLMs in security assertion generation. They perform formal property verification to identify vulnerabilities before developing the Design Under Test (DUT).
@@
In contrast, we have exclusively used the design specification in natural language to generate SVA instead of directly providing RTL code for assertion generation to the LLM. Since the objective of assertions is to identify the bug in the implementation, it is not desirable to obtain the assertions from implementation. Moreover, creating assertions from a design specification poses a significant challenge in the absence of a dedicated and specialized framework. We have also observed that the recent works are non-iterative, i.e., once the assertions are generated, the frameworks don’t try to refine any errors, such as syntactic or simulation errors, generated by LLM. Therefore, we proposed a novel framework, ChIRAAG, that formats the input specification and gives it to the LLM and further reviews, analyzes, and mends the LLM-generated assertions, if required. We have mainly focused on generating functional assertions rather than generating security assertions. None of the above-discussed works generate assertions using only the specification.
@@ III ChIRAAG: Our SVA Generation Framework
The automatic assertion generation with ChIRAAG involves two stages: (i) formatting of specifications and (ii) automatic assertion generation. Algorithm 1 shows the steps of our proposed ChIRAAG framework. The stages are briefly described below:
@@ III-A Formatting of specification
The formatting of design specifications includes representing the design context with specific labels, informing about its general definitions, signals, functional parameters, etc. The formatting involves extracting the important information from the design specification, $\mathcal{D}$, and later converting the extracted information, $UnFmt$, into a Javascript Object Notation (JSON) format, $\mathcal{J}$, using the functions Extraction and JSON, as mentioned in Line 6 of Algorithm 1. The $JSON$ function comprises several sub-functions as shown in Table I to produce the formatted specification.
@@ III-B Automatic Assertion Generation
In the next step, ChIRAAG generates SVA using the JSON formatted specifications. The formatted specifications of the designs are provided as input to the LLM through our framework by making a call to the OpenAI Application Programming Interface (API). This generates raw SVA as stated in Line 10 of Algorithm 1. The generated assertions are further verified for correctness in a simulation environment using testcases. This generates simulation logs that are then parsed using the ‘log_parser’ function, producing log message, $\mathcal{W}_{log}$, as shown in Line 13 of Algorithm 1. This step gives rise to four possible cases that can eventually occur. They are
@@
(a) No Error Message: When the generated assertion, $A_{n}$, passes all the test cases, i.e., no error message in $\mathcal{W}_{log}$, then $A_{n}$ is taken as correct output assertions suite.
@@
(b) Syntax Error: Some LLM-generated assertions have Syntax Errors (SE). The LLM is provided the log message, $\mathcal{W}_{log}$, automatically. Further, the LLM is prompted to refine the assertion. LLM is able to correct the SE within a few iterations. Line 17 of Algorithm 1 refine the assertion based on the error message.
(c) Simulation Failure: The assertions generated by LLM might fail in the simulation run because of specific errors. Two such simulation errors we have encountered are timing errors and missing signals. Line 20 of Algorithm 1 refine the assertion based on the error message.
@@
(d) Implementation Bug: It might be possible that even the generated assertions are correct, but the implementation might be buggy. In such a scenario, the testcase fails, giving wrong outputs. In this case, manual inspection of design implementation is to be done as shown in Line 23 of Algorithm 1.
@@
The assertions may not pass the testcases in a single run. Therefore, we iteratively run the error-checking process multiple times. Line 11 of Algorithm 1 shows that the while loop runs until $n$ is less than the upper bound $\mathcal{T}$. If $n$ gets equal to $\mathcal{T}$, and there are still error messages in $\mathcal{W}_{log}$, we manually investigate and resolve the potential issue which LLM cannot fix automatically and restart the assertion generation process as stated in Lines 29-33.
The proposed ChIRAAG framework is illustrated in the Fig. 1. We use the design only to check the syntax and semantics of the generated assertions. No other refinement of assertions is done using the design. In this process, a correctly generated assertion can fix a bug (if any) in a design as well.
@@
Algorithm 1 ChIRAAG: SVA Generation Framework
@@
2:      $\mathcal{D}$: Design specification document, $\mathcal{L}$: Log file of simulation tool, $\mathcal{T}$: Upper bound on the number of LLM interaction prompts
3:Output:
4:      $A_{n}$: Suite of System Verilog Assertions
5:procedure Specification Extraction and Formatting
6:    $UnFmt\leftarrow Extraction(\mathcal{D})$, $\mathcal{J}\leftarrow\text{JSON}(UnFmt)$
7:end procedure
8:procedure LLM Interaction for SVA Generation
9:    Initialize $n\leftarrow 0\text{ and }\mathcal{L}\leftarrow\{\}$
10:    $A_{n}\leftarrow\text{LLM}(\mathcal{J},\mathcal{L})$
11:    while $n<\mathcal{T}$ do
12:         Use VCS to check the generated $A_{n}$ on testcases
13:         $\mathcal{W}_{log}\leftarrow\text{log\_parser}(\mathcal{L})$
14:         if no errors in $\mathcal{W}_{log}$ then
15:             break
16:         else if syntactic errors in $\mathcal{W}_{log}$ then
17:             $\overline{A}_{n}\leftarrow$ Rectify syntactic errors
18:             $A_{n}\leftarrow\overline{A_{n}}$
19:         else if timing or missing signal error in $\mathcal{W}_{log}$ then
20:             $\overline{A}_{n}\leftarrow$ Rectify simulation errors
21:             $A_{n}\leftarrow\overline{A_{n}}$
22:         else if testcase failure message in $\mathcal{W}_{log}$ then
23:             Manual inspection of design implementation for bug
24:             break
25:         end if
26:         $n\leftarrow n+1$
27:    end while
28:end procedure
29:procedure Design Implementation Examination
30:    if $n==\mathcal{T}$ and error message in $\mathcal{W}_{log}$ then
31:         Examine design implementation and restart the process
32:    end if
33:end procedure
@@ IV Case Study: RV Timer
RV Timer [5] is a design module that is intended for use by processors to monitor the current time relative to the system reset or power-on. The block diagram of the RV Timer is shown in Fig. 2
@@
The unstructured design specification of RV Timer is initially transformed to structured JSON format and presented as input to the LLM to generate the SVA. The LLM initially generated ten functional assertions. We further evaluated the correctness of the raw assertions using testcases and VCS simulator. Seven out of ten assertions were correct (i.e., there are no syntax or simulation errors) in their raw format. This shows the potential of our proposed framework ChIRAAG. The remaining three raw assertions were rectified automatically using the steps of Algorithm 1. This process generated an extra assertion, which satisfies the design aspects, increasing the count of generated assertions to eleven. Some of the initial assertions of RV Timer generated by the LLM are:
@@
The Assertion 1 is a correct raw assertion that passes all the tests. It ensures that ‘tick_count’ resets to 0 when reset is asserted, i.e., when ‘rst_ni’ or the active low reset signal is not active. The Assertion 2, we observed that a missing signal prevents passing all the test cases. The LLM understood the missing signal rst_ni from the simulation log message and included it in the generated assertion in the next iteration. Line 20 of Algorithm 1 is called upon to rectify the Assertion 2. The LLM generated the refined assertion and modified it into two distinct assertions Modified Assertion 2a and Modified Assertion 2b.
@@
The Modified Assertion 2a ensures that as long as the system is not in reset, is active, and the tick count is below a set threshold, the tick count will increment by 1 in each clock cycle. The Modified Assertion 2b verifies that when the system transitions from a reset state to a non-reset state, the tick count is reset to 0, ensuring proper initialization or re-initialization of the system’s tick counter. ChIRAAG took twelve iterations to converge because, initially, it has syntax errors like identifier missing and token missing, which are resolved in seven iterations for all the assertions combined, and then simulation errors like timing failure are rectified using further prompting in five more iterations.